OpenStudy (anonymous):

Anyone help? Using complete sentences, compare and contrast the ways to factor the "sum of two squares" and the "difference of two squares." Provide an explanation and example of each, including any similarities and differences.

OpenStudy (helder_edwin):

first of all u cannot factor the sum of two squares, unless u are allowed to use complex numbers, in such case \[ \large x^2+y^2=(x+yi)(x-yi) \] secon the difference of two squares can be factored into \[ \large x^2-y^2=(x+y)(x-y) \]

OpenStudy (anonymous):

Thanks c:

OpenStudy (helder_edwin):

u r welcome